# Rua Dr. Bento Teobaldo Ferraz 271, 01140-070, São Paulo, SP, Brazil

### v2

@@ Abstract
Coupling $N$ large rank $m$ minimal models and flowing to IR fixed points is a systematic way to build new classes of compact unitary 2d CFTs which are likely to be irrational, and potentially have a positive Virasoro twist gap above the vaccuum. In this paper, we build on the construction of Antunes:2022vtb , establishing that, for spins less than 10, additional currents transforming in non-trivial irreducible representations of the permutation symmetry $S_{N}$ are not conserved at the IR fixed points. Along the way, we develop a finer understanding of the spectrum of these theories, of the special properties of the $N=4$ case and of non-invertible symmetries that constrain them. We also discuss variations of the original setup of Antunes:2022vtb , some of which can exist for smaller values of the UV central charge.
@@ 1 Introduction
As discussed in grz18 , $\epsilon^{i}$ has holomorphic dimension $\frac{3}{f}-\frac{1}{2}$ in the ultraviolet (UV) where $q=2+2\cos(\pi f/2)$ and $2\leq f\leq 4$. Although the UV CFT in (1) is solvable and unitary for $q=3$, the renormalization group (RG) flow away from it is strongly coupled. To make it weakly coupled, we must give up unitarity by taking $q$ to be close to $2$ or $4$.
The construction of Antunes:2022vtb extends the range of potential unitary infrared CFTs by coupling higher rank $m$ minimal models and using $1/m$ as an expansion parameter, leading to infinite sequences of unitary weakly coupled fixed points as $1/m$ approaches 00. A key property of these fixed points is that additional $S_{N}$ singlet currents (checked up to spin $J\leq 10$) of the UV tensor product theory stop being conserved in the infrared. This is evidence for irrationality due to modular invariance and the fact that $c>1$ for these models. In this work, we extend the results of Antunes:2022vtb in several directions:
@@
    When the symmetry is ungauged, we classify the currents in all non-trivial representations of $S_{N}$ and show that they acquire anomalous dimensions. We further argue that this should remain true at large but finite $m$.
@@
    We show that the $N=4$ system preserves a large class of non-invertible symmetries and give a non-perturbative argument for the integrability of an associated deformation.
@@
This paper is organized as follows. In section 2, we review the $S_{N}$ symmetric models of Antunes:2022vtb and the technique that was used to decide between conservation and non-conservation of the singlet UV currents. In particular, we advocate for a more thorough check which goes beyond the $S_{N}$ singlet sector. Next, in section 3, we sharpen our expectations regarding the non-singlet sector by combining standard 2d CFT techniques with $S_{N}$ representation theory. Section 4 then uses this understanding to develop algorithms which are able to perform the desired more powerful check. This leads to the result that there are still no signs of any chiral symmetry beyond Virasoro. Extensions of the models studied here along with an analysis of non-invertible symmetries in both the original and extended models are discussed in section 5 before we give an outlook on future directions in section 6.
@@ 2.1 The model
The models considered in Antunes:2022vtb , which this paper will study in more detail, have been designed to circumvent the problem with the left diagram of Figure 1 in favour of the right diagram. There are many tensor products of solvable CFTs which allow one to construct weakly coupled RG flows to less familiar fixed points. But simply constructing flows which become arbitrarily short is not enough. The theories from which they emanate must also be unitary if one is to have a hope of reaching the physically interesting case of a compact unitary irrational CFT with minimal chiral symmetry. This makes it quite natural to take rank $m$ minimal models as our seed theories and work in the regime where unitary minimal models accumulate. This gives us access to large-$m$ perturbation theory which was first used in Zamolodchikov:1987ti ; lc87 and later in dns01 ; Behan:2021tcn for different purposes.
@@
$\begin{split}S_{\text{CMM}}=\sum_{i=1}^{N}S_{m}^{i}+\int d^{2}x\,g_{\sigma}\sigma+g_{\epsilon}\epsilon,\end{split}$ (8)
@@
up to $O(g^{3})$ terms. Setting (13) to zero leads to four fixed points. The obvious ones are $(g_{\sigma}^{*},g_{\epsilon}^{*})=(0,0)$ which is fully unstable and $(g_{\sigma}^{*},g_{\epsilon}^{*})=\left(0,\frac{2\sqrt{3}}{m\pi}\right)$ which is fully stable (following immediately from the fact that it is a decoupled fixed point). [^10] Stable and unstable here are being used in the RG sense. It is expected that all coupled minimal model fixed points have a stable vaccuum because they are not perturbations around a free theory. The latter is nothing but $N$ minimal models of rank $m-1$ Zamolodchikov:1987ti ; lc87 . The more interesting fixed points $\text{FP}^{*}_{\pm}$ are
@@
Figure 2: Schematic diagram of the fixed points reproduced from Antunes:2022vtb . The tricritical fixed points in the middle given by (14) are the conjectured examples of compact unitary irrational CFTs with only Virasoro symmetry.
@@ 2.2 Anomalous dimensions of singlet currents
To find evidence of irrationality which is actually convincing, Antunes:2022vtb undertook a search for generators of an extended chiral algebra in the IR and found none. Due to its large $\mathfrak{Vir}^{N}$ chiral algebra, the UV has many such currents but none of them were found to still be conserved in the IR. Specifically, this check was done for spin $J\leq 10$ currents in the singlet representation of $S_{N}$ which transform as primaries under the diagonal $\widehat{\mathfrak{Vir}}$ generated by the sum of stress tensors. Checking primaries is sufficient because $\widehat{\mathfrak{Vir}}$ Verma modules have to lift together but the singlet assumption will be relaxed very soon.
When a UV current fails to be conserved in the IR, this means that it acquires an anomalous dimension. Computing this anomalous dimension directly requires conformal perturbation theory at two loops. Indeed, $T^{K}\mathcal{O}T^{L}\right>=0$ whenever $T^{K}$ is a current and $\mathcal{O}\in\{\sigma,\epsilon\}$ is a deforming operator. Fortunately, Antunes:2022vtb was able to avoid a two-loop calculation by using a more modern method which exploits the non-conservation equation
@@
The idea is that when the operator content of the theory changes continuously, taking $g_{\sigma}\to 0$ in (16) leaves behind a short multiplet with $(h,\bar{h})=(J,0)$ and a long multiplet with $(h,\bar{h})=(J,1)$. These must recombine when the coupling is turned on. We have allowed $b^{k}_{\;\;L}\neq\delta^{K}_{L}$ in (16) because in practice we must start by generating a basis of candidates for the divergence of $T^{K}$ with the precise linear combination being found later. Once $b^{K}_{\;\;L}$ is known, the anomalous dimension of the broken current follows from
@@
$\begin{split}g_{\sigma}\int d^{2}z_{3}\,\left<\bar{\partial}T^{K}(z_{1})V^{L}(z_{2},\bar{z}_{2})\sigma(z_{3},\bar{z}_{3})\right>.\end{split}$ (18)
On the other hand, (16) turns the same correlator into
@@
Setting (18) equal to (19), we can solve for $b^{K}_{\;\;L}$ by integrating something which localizes to a delta function. A key fact is that the only way for $T^{K}$ to stay conserved is to have $b^{K}_{\;\;L}=0$ as a matrix. This can only happen if all of the three-point functions in (18) vanish.
While $b^{K}_{\;\;L}\neq 0$ for a single $L$ is sufficient to conclude that $T^{K}$ is not a dilation eigenstate with weights of $(h,\bar{h})=(J,0)$ in the IR, we are more interested in establishing this for arbitrary linear combinations of the $T^{K}$. As such, the check to undertake is that $b^{K}_{\;\;L}$ is a matrix whose rank is equal to the number of rows. This is what led to the conclusion in Antunes:2022vtb that $S_{N}$ singlets up to spin-10 for $N=5,6,7$ all lift. The next two sections are about applying the same strategy to non-singlets.
@@ 2.3 When are singlets enough?
due to dmvv96 which computes the modular invariant partition function of any symmetric product orbifold in terms of the modular invariant partition function of the seed theory and so called Hecke operators.
Before developing technology for non-singlet currents, which will be important for the non-orbifold theory, let us point out that conformal perturbation theory for symmetric orbifolds has recently come to play a large role in studies of $AdS_{3}/CFT_{2}$. In the most well established example egg18 ; egg19 , the seed theory is a sigma model for $K3$ or $T^{4}$ and the subsequent deformation is by an exactly marginal operator in the twisted sector. In this case, higher-spin currents (and other short multiplets) are again lifted but the deformed theories have more than just Virasoro symmetry since they are supersymmetric. Some work in this direction, including the identification of Regge trajectories, has been done in gpz15 ; bkz21 . A less supersymmetric version of $AdS_{3}/CFT_{2}$ using $\mathcal{N}=2$ minimal models, which has been proposed in bbchk20 ; abbck22 ; bbch22 , also exhibits lifting of higher-spin currents as one moves toward irrational SCFTs on the moduli space. A possible inverse to this phenomenon, wherein one encoutners new rational points, has been studied perturbatively in bkoz20 ; k23 .
@@ 3 Permutation symmetry and charged currents
The technique in section 2.2, which alleviates the need for direct two-loop conformal perturbation theory, makes it crucial to understand the spaces of both UV currents and the operators with the right quantum numbers to recombine with them. Virasoro characters provide an efficient starting point for this task. In particular, they allow us to obtain the dimension of either space for arbitrarily high spins and to refine this dimension by $S_{N}$ representation for certain low spins. This section will explain the steps involved focusing mostly, but not exclusively, on the large $m$ regime where perturbative computations are most reliable.
@@ 3.2 Counting currents and divergences
$\chi^{(c)}_{h}(q)=\text{Tr}_{\mathcal{V}_{c,h}}(q^{L_{0}-\frac{c}{24}})\,,$ (30)
@@
$\chi_{(1,1)}^{(c\to 1)}(q)=\frac{q^{-\frac{1}{24}}}{\phi(q)}(1-q)\,,\quad\chi_{(r,s)}^{(c\to 1)}(q)=\frac{q^{h_{r,s}-\frac{1}{24}}}{\phi(q)}(1-q^{rs})\,,$ (31)
@@
$Z_{m\to\infty}(q)=\sum_{1\leq r\leq s}^{\infty}|\chi_{(r,s)}^{(c\to 1)}(q)|^{2}\,.$ (32)
@@
$\begin{split}Z_{\textrm{CMM, UV}}&=\sum_{h,\bar{h}}d_{h,\bar{h}}\,\chi_{h}^{(c=N)}(q)\chi_{\bar{h}}^{(c=N)}(\bar{q})\\ &\supset\sum_{n=0}^{\infty}d_{n,0}\chi_{n}^{(c=N)}(q)(1+O(\bar{q}))+d_{n,1}\chi_{n}^{(c=N)}(q)(\bar{q}+O(\bar{q}^{2}))\,,\end{split}$ (33)
@@
$\chi^{(N)}_{h}(q)=\frac{q^{h-\frac{N}{24}}}{\phi(q)}\,,\quad\phi(q)=\prod_{k=1}^{\infty}(1-q^{k})\,,$ (34)
@@
$\begin{split}\sum_{n=0}^{\infty}d_{n,0}\chi_{n}^{(c=N)}(q)&=\left(\chi_{(1,1)}^{(c\to 1)}(q)\right)^{N}\,,\\ \quad\sum_{n=0}^{\infty}d^{\prime}_{n,1}\chi_{n}^{(c=N)}(q)&=\binom{N}{4}\left(\chi_{(1,1)}^{(c\to 1)}(q)\right)^{N-4}\left(\chi_{(1,2)}^{(c\to 1)}(q)\right)^{4}\,,\end{split}$ (35)
where the prime in the second sum denotes that we selected only operators that contain four $\phi_{(1,2)}$ constituents, since only these are relevant for the subsequent computations of anomalous dimensions. Expanding both sides at small $q$ allows us to determine the degeneracies to an arbitrarily high order, but we were not able to guess a closed form expression. We therefore list the degeneracies up to spin 10 for arbitrary $N$ in Table 2.

### v3

@@ Abstract
Coupling $N$ large $m$ minimal models and flowing to IR fixed points is a systematic way to build new classes of compact unitary 2d CFTs which are likely to be irrational, and potentially have a positive Virasoro twist gap above the vacuum. In this paper, we build on the construction of Antunes:2022vtb , establishing that, for spins less than 10, additional currents transforming in non-trivial irreducible representations of the permutation symmetry $S_{N}$ are not conserved at the IR fixed points. Along the way, we develop a finer understanding of the spectrum of these theories, of the special properties of the $N=4$ case and of non-invertible symmetries that constrain them. We also discuss variations of the original setup of Antunes:2022vtb , some of which can exist for smaller values of the UV central charge.
@@ 1 Introduction
As discussed in grz18 , CFT data in ultraviolet (UV) theory is conveniently specified in terms of a coupling constant $f$ such that $q=2+2\cos(\pi f/2)$. The critical and tricritical $q$-state Potts models, which annihilate at $q=4$, are described in the range $q\in[0,4]$ by the branches $f\in[2,4]$ and $f\in[4,6]$ respectively. In this notation, $\epsilon^{i}$ starts off with the holomorphic dimension $\frac{3}{f}-\frac{1}{2}$ which means $\epsilon^{i}\epsilon^{j}$ is only classically marginal for $q=2$. Although the tensor product CFT in (1) is solvable and unitary for $q=3$, the renormalization group (RG) flow away from it is strongly coupled. To make it weakly coupled, we must give up unitarity by taking $q$ to be close to $2$.
The construction of Antunes:2022vtb extends the range of potential unitary infrared CFTs by instead coupling minimal models indexed by $m$ and using $1/m$ as an expansion parameter, leading to infinite sequences of unitary weakly coupled fixed points as $1/m$ approaches 00.555In the convention we use, the first unitary minimal model (the critical Ising model) corresponds to $m=3$. A key property of these fixed points is that additional $S_{N}$ singlet currents of the UV tensor product theory (checked up to spin $J\leq 10$) stop being conserved in the infrared. Evidence for the breaking of chiral symmetry is, at the same time, evidence for irrationality because these models are modular invariant with $c>1$. The growth estimates in Cardy:1986ie ; Mukhametzhanov:2019pzy , showing that there are infinitely many Virasoro primaries, are therefore applicable meaning that an RCFT is not possible without an extended chiral algebra. In this work, we extend the results of Antunes:2022vtb in several directions:
@@
    When the symmetry is ungauged, we classify the currents in all non-trivial representations of $S_{N}$ and show that those below spin 10 acquire anomalous dimensions. We further argue that this should remain true at large but finite $m$.
@@
    For the $N=4$ system, we show that it preserves a large class of non-invertible symmetries and give a non-perturbative argument for the integrability of an associated deformation.
@@
This paper is organized as follows. In section 2, we review the $S_{N}$ symmetric models of Antunes:2022vtb and the technique that was used to decide between conservation and non-conservation of the singlet UV currents. In particular, we advocate for a more thorough check which goes beyond the $S_{N}$ singlet sector. Next, in section 3, we sharpen our expectations regarding the non-singlet sector with the use of torus partition functions and $S_{N}$ representation theory. Section 4 then uses this understanding to develop algorithms which are able to perform the desired more powerful check. This leads to the result that there are still no signs of any chiral symmetry beyond Virasoro. Extensions of the models studied here along with an analysis of non-invertible symmetries in both the original and extended models are discussed in section 5 before we give an outlook on future directions in section 6.
@@ 2.1 The model
The models considered in Antunes:2022vtb , which this paper will study in more detail, have been designed to circumvent the problem with the left diagram of Figure 1 in favour of the right diagram. There are many tensor products of solvable CFTs which allow one to construct weakly coupled RG flows to less familiar fixed points. But simply constructing flows which become arbitrarily short is not enough. The theories from which they emanate must also be unitary if one is to have a hope of reaching the physically interesting case of a compact unitary irrational CFT with minimal chiral symmetry. This makes it quite natural to take minimal models as our seed theories and work in the regime where unitary minimal models accumulate, namely $m\to\infty$. This gives us access to large-$m$ perturbation theory which was first used in Zamolodchikov:1987ti ; lc87 and later in dns01 ; Behan:2021tcn for different purposes.
@@
$\begin{split}S_{\text{CMM}}=\sum_{i=1}^{N}S_{m}^{i}+\int d^{2}x\,(g_{\sigma}\sigma+g_{\epsilon}\epsilon),\end{split}$ (8)
@@
up to $O(g^{3})$ terms. Setting (13) to zero leads to four fixed points. The obvious ones are $(g_{\sigma}^{*},g_{\epsilon}^{*})=(0,0)$ which is fully unstable and $(g_{\sigma}^{*},g_{\epsilon}^{*})=\left(0,\frac{\sqrt{3N}}{m\pi}\right)$ which is fully stable (following immediately from the fact that it is a decoupled fixed point). [^11] Stable and unstable here are being used in the RG sense. It is expected that all coupled minimal model fixed points have a stable vacuum because they are not perturbations around a free theory. The latter is nothing but $N$ copies of the minimal model with $m$ replaced by $m-1$ Zamolodchikov:1987ti ; lc87 . The more interesting fixed points $\text{FP}^{*}_{\pm}$ are
@@
Figure 2: Schematic diagram of the fixed points reproduced from Antunes:2022vtb . The tricritical fixed points $\text{FP}^{*}_{\pm}$ given by (14) are the conjectured examples of compact unitary irrational CFTs with only Virasoro symmetry.
@@ 2.2 Anomalous dimensions of singlet currents
To find evidence of irrationality which is actually convincing, Antunes:2022vtb undertook a search for generators of an extended chiral algebra in the IR and found none. Due to its large $\mathfrak{Vir}^{N}$ chiral algebra, the UV has many such currents but none of them were found to still be conserved in the IR. Specifically, this check was done for spin $J\leq 10$ currents in the singlet representation of $S_{N}$ which transform as primaries under the diagonal $\widehat{\mathfrak{Vir}}$ generated by the sum of stress tensors. Since $\widehat{\mathfrak{Vir}}$ remains unbroken in the IR, local operators must arrange themselves into representations with respect to this symmetry. Within a given multiplet, the anomalous dimensions (if any) are the same for the primaries and descendants so it is sufficient to only check the former. On the other hand, the singlet assumption will be relaxed very soon.
When a UV current transforming as a primary under $\widehat{\mathfrak{Vir}}$ fails to be conserved in the IR, this means that its anomalous dimension is non-zero. Computing this anomalous dimension directly requires conformal perturbation theory at two loops. Indeed, $T^{K}\mathcal{O}T^{L}\right>=0$ whenever $T^{K}$ is a spin-$J$ current and $\mathcal{O}\in\{\sigma,\epsilon\}$ is a deforming operator. Fortunately, Antunes:2022vtb was able to avoid a two-loop calculation by using a more modern method which exploits the non-conservation equation
@@
The initial observation is that when the operator content of the theory changes continuously, taking $g_{\sigma}\to 0$ in (16) leaves behind a short multiplet with $(h,\bar{h})=(J,0)$ and a long multiplet with $(h,\bar{h})=(J,1)$. For a given spin $J$, it is conceptually straightforward to enumerate a basis of $(h,\bar{h})=(J,0)$ operators whose elements will be denoted $T^{K}$. These are what we have been referring to as (UV) currents. It is similarly possible to produce a basis for the space of $(h,\bar{h})=(J,1)$ operators with elements $V^{K}$. In practice, one should regard these bases as arbitrary which means they are almost certainly not nice enough to yield $T^{K}$ and $V^{K}$ operators which match up one-to-one. As such, the divergence of a given $T^{K}$ at one loop is a non-trivial linear combination of the $V^{K}$ whose coefficients are encoded in the all-important matrix $b^{K}_{\;\;L}$. Before this linear combination is taken, the $V^{K}$ will be referred to as candidates for being the divergence of a spin-$J$ current. We will also employ the phrase potential divergences, dropping the word “potential” for brevity when the context makes this clear. [^12] Since we are studying the lifting induced by $\sigma$, one more condition we will often impose is that four copies of $\phi_{(1,2)}$ must be present in the expression for $V^{K}$ to make it a divergence candidate or a potential divergence.
The multiplets described above must recombine when the coupling $g_{\sigma}$ is turned on and this leads to a concrete algorithm for determining which currents lift. This revolves around computing $b^{K}_{\;\;L}$ which yields the precise linear combination of divergence candidates that enters in (16). It also fixes the anomalous dimension $\gamma$ of the broken current since
@@
$\begin{split}g_{\sigma}\int d^{2}z_{3}\,\left<\bar{\partial}T^{K}(z_{1})V^{L}(z_{2},\bar{z}_{2})\sigma(z_{3},\bar{z}_{3})\right>\end{split}$ (18)
where the integrand is a three-point function in the undeformed theory. On the other hand, (16) turns the same correlator into
@@
Setting (18) equal to (19), we can solve for $b^{K}_{\;\;L}$ by performing a calculation at one-loop instead of two. Moreover, the three-point function at zero loops that we are instructed to integrate vanishes at separated points due to ¯$\bar{\partial}T^{K}(z_{1})$. This leads to the appearance of a delta function which makes the integral trivial.
Due to the simplification of this integral, one can observe a key fact which is that the only way for $T^{K}$ to stay conserved is to have $b^{K}_{\;\;L}=0$ for all $L$. This can only happen if the three-point functions in (18) similarly vanish for all $L$. While $b^{K}_{\;\;L}\neq 0$ for a single $L$ is sufficient to conclude that $T^{K}$ is not a dilation eigenstate with weights of $(h,\bar{h})=(J,0)$ in the IR, we are more interested in establishing this for arbitrary linear combinations of the $T^{K}$. As such, the check to undertake is that $b^{K}_{\;\;L}$ is a matrix whose rank is equal to the number of rows. This is what led to the conclusion in Antunes:2022vtb that $S_{N}$ singlets up to spin-10 for $N=5,6,7$ all lift. The next two sections are about applying the same strategy to non-singlets.
@@ 2.3 When are singlets enough?
due to dmvv96 which computes the partition function of any symmetric product orbifold using Hecke operators which act on the partition function of the seed theory according to
$\begin{split}T_{k}Z(\tau,\bar{\tau})=\sum_{d|k,a=\frac{k}{d}}\sum_{b=0}^{d-1}Z\left(\frac{a\tau+b}{d},\frac{a\bar{\tau}+b}{d}\right).\end{split}$ (22)
The defining property of Hecke operators now ensures that modular invariance of $Z_{S_{N}}(\tau,\bar{\tau})$ follows immediately from modular invariance of $Z(\tau,\bar{\tau})$. [^13] To see that $T_{k}Z(\tau,\bar{\tau})$ is modular invariant, note that (22) is a sum over upper triangular matrices $M$ with determinant $k$. When computing $T_{k}Z(\gamma\tau,\gamma,\bar{\tau})$ for $\gamma\in SL(2,\mathbb{Z})$, $M\gamma$ is no longer upper triangular but it can be brought back to this form by acting with another $SL(2,\mathbb{Z})$ matrix on the left. This preserves $Z(\tau,\bar{\tau})$ by assumption. The orbifold by $S_{N}$ is therefore well defined for any tensor product CFT.
Before developing technology for non-singlet currents, which will be important for the non-orbifold theory, let us point out that conformal perturbation theory for symmetric orbifolds has recently come to play a large role in studies of $AdS_{3}/CFT_{2}$. In the most well established example egg18 ; egg19 , the seed theory is a sigma model for $K3$ or $T^{4}$ and the subsequent deformation is by an exactly marginal operator in the twisted sector. In this case, higher-spin currents (and other short multiplets) are again lifted but the deformed theories have more than just Virasoro symmetry since they are supersymmetric. Some work in this direction, including the identification of Regge trajectories, has been done in gpz15 ; bkz21 . A less supersymmetric version of $AdS_{3}/CFT_{2}$ using $\mathcal{N}=2$ minimal models, which has been proposed in bbchk20 ; abbck22 ; bbch22 , also exhibits lifting of higher-spin currents as one moves toward irrational SCFTs on the moduli space. A possible inverse to this phenomenon, wherein one encounters new rational points, has been studied perturbatively in bkoz20 ; k23 .
@@ 3 Permutation symmetry and charged currents
The technique in section 2.2, which alleviates the need for direct two-loop conformal perturbation theory, makes it crucial to understand the spaces of both UV currents and the operators with the right quantum numbers to recombine with them. Virasoro characters provide an efficient starting point for this task. In particular, they allow us to obtain the dimension of either space for arbitrarily high spins and to refine this dimension by $S_{N}$ representations. This section will explain the steps involved focusing mostly, but not exclusively, on the large $m$ regime where perturbative computations are most reliable.
@@ 3.2 Counting currents and divergences
$\chi^{(c)}_{h}(\tau)=\text{Tr}_{\mathcal{V}_{c,h}}(q^{L_{0}-\frac{c}{24}})\,,\quad q\equiv\exp(2\pi i\tau)\,,$ (31)
@@
$\chi_{(1,1)}^{(c\to 1)}(\tau)=\frac{q^{-\frac{1}{24}}}{\phi(q)}(1-q)\,,\quad\chi_{(r,s)}^{(c\to 1)}(\tau)=\frac{q^{h_{r,s}-\frac{1}{24}}}{\phi(q)}(1-q^{rs})\,,$ (32)
@@
$Z_{m\to\infty}(\tau,\bar{\tau})=\sum_{1\leq r\leq s}^{\infty}\left|\chi_{(r,s)}^{(c\to 1)}(\tau)\right|^{2}\,.$ (33)
@@
$\begin{split}Z_{\textrm{CMM, UV}}(\tau,\bar{\tau})&=\sum_{h,\bar{h}}d_{h,\bar{h}}\,\chi_{h}^{(c=N)}(\tau)\chi_{\bar{h}}^{(c=N)}(\bar{\tau})\\ &\supset\sum_{n=0}^{\infty}d_{n,0}\chi_{n}^{(c=N)}(\tau)(1+O(\bar{q}))+d_{n,1}\chi_{n}^{(c=N)}(\tau)(\bar{q}+O(\bar{q}^{2}))\,,\end{split}$ (34)
@@
$\chi^{(N)}_{h}(\tau)=\frac{q^{h-\frac{N}{24}}}{\phi(q)}\,,\quad\phi(q)=\prod_{k=1}^{\infty}(1-q^{k})\,,$ (35)
@@
$\begin{split}\sum_{n=0}^{\infty}d_{n,0}\chi_{n}^{(c=N)}(\tau)&=\left(\chi_{(1,1)}^{(c\to 1)}(\tau)\right)^{N}\,,\\ \quad\sum_{n=0}^{\infty}d^{\prime}_{n,1}\chi_{n}^{(c=N)}(\tau)&=\binom{N}{4}\left(\chi_{(1,1)}^{(c\to 1)}(\tau)\right)^{N-4}\left(\chi_{(1,2)}^{(c\to 1)}(\tau)\right)^{4}\,,\end{split}$ (36)
where the prime in the second sum denotes that we selected only operators that contain four $\phi_{(1,2)}$ constituents, since only these are relevant for the subsequent computations of anomalous dimensions. Expanding both sides at small $q$ allows us to determine the degeneracies to an arbitrarily high order, but we were not able to guess a closed form expression. We therefore list the degeneracies up to spin 10 for arbitrary $N$ in Table 2. These degeneracies turn out to be a polynomial in $N$ of degree $\lfloor J/2\rfloor$.